OpenStudy (anonymous):

how 3 x 6 + 2x0 = a.6 b.20 c.18 d.22

OpenStudy (anonymous):

raes001, I can't understand the format of your equation is it: \[3 \times 6 + 2 \times 0\]

OpenStudy (anonymous):

how 1+1+1+1+1+1x0

OpenStudy (aripotta):

wait, is that a whole new question?

OpenStudy (aripotta):

and anything multiplied by 0 is 0

OpenStudy (anonymous):

The last one is just 5.. you don't make any sense?

OpenStudy (aripotta):

:l

OpenStudy (anonymous):

It's 1+1+1+1+1+(1*0) = 5

OpenStudy (aripotta):

yea, i got that after i posted lol

OpenStudy (anonymous):

P Parentheses first E Exponents (ie Powers and Square Roots, etc.) MD Multiplication and Division (left-to-right) AS Addition and Subtraction (left-to-right)

OpenStudy (anonymous):

However due to the priority rules the first question is (3*6) + (2*0) = 18

OpenStudy (anonymous):

This would mean that frx's rewrite of the equation is correct for your second equation. However, if the format of the first equation is how I wrote it, then there wasn't a correct answer in the options offered.

OpenStudy (anonymous):

yes, frx and I are in agreement. That is what I would get, too, using PEMDAS.

OpenStudy (anonymous):

any question for me?